### v2

# Sensing with near-infrared laser trapped fluorescent nanodiamonds

Abstract
Biosensing based on optically trapped fluorescent nanodiamonds is an intriguing research direction potentially allowing to resolve biochemical processes inside living cells. Towards this goal, we investigate near infrared (NIR) laser irradiation at 1064 nm on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers. The NIR wavelength is a popular choice for optical trapping in biological samples because of its low absorption in bio-samples. By conducting comprehensive experiments, we aim to understand how NIR exposure influences the fluorescence and sensing properties of FNDs and to determine the potential implications for the use of FNDs in various sensing applications. The experimental setup involved exposing FNDs to varying intensities of NIR laser light and analyzing the resultant changes in their optical and physical properties. Key measurements included all-optical fluorescence relaxation, optical spectroscopy, and optically detected magnetic resonance (ODMR) spectra. The findings reveal how increased NIR laser power correlates with alterations in ODMR central frequency but also that charge state dynamics under NIR irradiation of NV centers plays a role. We suggest protocols with NIR and green light that mitigate the effect of NIR, and demonstrate that FND biosensing works well with such a protocol.
Nanodiamond, Optical trapping, Temperature sensing, NV-centers
preprint: APS/123-QED
I Introduction
Figure 1: Graphical abstract illustrating the sensing modalities explored to investigate effects of NIR laser on NV center for sensing. Temperature, abundance of paramagnetic species, and pH, are altered while optical spectroscopy, ODMR spectra, and fluorescence polarization relaxometry is measured from optically trapped nanodiamonds.
Fluorescent nanodiamonds (FNDs) hosting negatively-charged nitrogen-vacancy (NV-) centers combine bright, photostable fluorescence with quantum-coherent electron-spin states, making them uniquely suited for bio-sensing and bio-imaging at the nanoscale [Schirhagl2014, Liu2021]. Because NV spins can be optically polarised and read out at room temperature, changes in their spin-state populations translate directly into optical signals, enabling quantitative sensing of magnetic fields, temperature and local electro-chemical potentials [Doherty2013, Rondin2014]. The ultimate performance of those sensors, however, is governed by the spin- and charge-state photodynamics of the NV centers themselves; understanding how an experimental control parameter perturbs those dynamics is therefore a prerequisite for accurate, high-sensitivity measurements.
The NV center is an electronic spin-1 system with a ground-state zero-field splitting (ZFS) $D\approx 2.87~\mathrm{GHz}$ that separates the $m_{s}=0$ and $m_{s}=\pm 1$ sub-levels. External magnetic fields lift the degeneracy of the $m_{s}=\pm 1$ levels through the Zeeman effect, while lattice strain, temperature and electric fields shift the ZFS and the optical transition energies via the Stark effect [Maze2008, Dolde2011]. In addition, the relative abundance of the neutral (NV0) and negative (NV-) charge states—and thus the fluorescence spectrum and intensity—depends sensitively on the local Fermi level and on photo-ionisation/recombination pathways. These intertwined dependencies let NV centers act as multi-modal nano-probes, provided that the separate interaction between all external controls (laser beams, microwave fields, etc.) and the FNDs are themselves well understood and do not compromise sensor performance.
A particularly powerful external control is optical trapping. Near-infra-red (NIR) optical tweezers can confine—and manoeuvre—sub-100  nm diamonds inside living cells or complex micro-fluidic environments, offering spatially resolved sensing with a resolution set by the nanoparticle size rather than by the optical diffraction limit [Neukirch2013]. Operating the trap at 1064  nm keeps tissue absorption low and mitigates photo-damage [Svoboda1994], yet the same intense NIR field may also heat the nanodiamond, alter NV charge-state kinetics or modify spin-lattice relaxation. Earlier work on bulk diamond or levitated nanodiamonds indeed reported substantial NIR-induced changes in fluorescence and $T_{\rm relax}$[Aslam2013, Ji2016], but a clear picture for FNDs dispersed in aqueous media—arguably the configuration of greatest practical relevance-is missing.
Here, as illustrated by Fig.1, we address that gap. Using freely diffusing and optically trapped FNDs in water and in biologically relevant buffer solutions, we quantify how a 1064  nm focused trapping beam (30–60  mW average power) affects NV optical and spin properties. We employ tree complementary read-outs:
  1. 1.
    photoluminescence (PL) spectroscopy to monitor charge-state redistribution [Geiselmann2013];
  2. 2.
    optically detected magnetic resonance (ODMR) to track shifts in the resonance frequencies with kHz precision [Gruber1997, Sow2020]; and
  3. 3.
    fluorescence relaxometry to probe magnetic-noise spectra—e.g. from free radicals or $\mathrm{Gd}^{3+}$ ions—at the NV transition frequency [Hall2016].
Together, these observables allow us to disentangle three mechanisms that NIR light can induce:
  1. 1.
    Photothermal heating of the FND, shifting the ZFS via the temperature coefficient $dD/dT\approx-74~\mathrm{kHz\,K^{-1}}$[Acosta2010];
  2. 2.
    Power-dependent ionisation/recombination that re-weights NV- and NV0, thereby changing both ODMR spectral properties and PL brightness [Aslam2013, Ji2016];
  3. 3.
    Modified surface sensitivity, whereby the NIR field—through either heating or altered charge state—amplifies or suppresses the NV response to environmental parameters such as pH or paramagnetic species concentration [Fu2010].
We first benchmark these effects in air-dried FNDs, then in trapped FNDs immersed in solutions of controlled pH and temperature, and finally in the presence of a paramagnetic relaxant ($\mathrm{Gd}^{3+}$). By comparing experiments with and without the NIR trap, we establish quantitative limits on NIR-induced artefacts and delineate operating regimes where accurate, spatially resolved biosensing is possible.
II Methods
II.1 Sample Preparation
Fluorescent FNDs with a nominal diameter of 120 nm (Sigma-Aldrich, 798088) were used for most of the experiments reported in this article if not stated otherwise. They possess a significant number of nitrogen-vacancy (NV) centers due to a high concentration of 3ppm. To prepare the FND suspension, 1 µL of the commercial stock solution of FNDs was diluted in 300-700 µL of deionized water depending on the desired concentration and before sonicating for 10 minutes. Next, the FND suspension was deposited on a petri dish (Mattek, part no P35G-1.5-10-C.), thoroughly cleaned to remove any contaminants that could interfere with the experiments and plasma ionised in order to make the surface hydrophilic, ensuring an even distribution of FNDs on the surface. Plasma treatment was applied for 30 seconds in a simple custom-device with a closed chamber in PMMA and under ambient conditions. Plasma was generated with a BD-20AV high-frequency generator (230 V, ETP, IL, USA) with an electrode positioned close to the petri dish surface. The plasma treatment modifies the surface energy of the petri dish by breaking chemical bonds on the surface and introducing polar functional groups, thereby inducing the hydrophilic state [Al-Azzam2022]. Approximately 3 µL of the FND suspension was then carefully deposited onto the clean petri dish using a micropipette. The suspension was gently spread across the surface before allowing it to dry. The drying process took place at room temperature to minimize any potential alterations to the FND properties potentially caused by excessive heating or cooling. After the solution had completely evaporated, the sample was ready for the optical experiments described in the following. The preparation of FNDs for experiments in the liquid environment also starts with the dilution steps and cleaning of the petri dish, but to maintain the FNDs in solution, the petri dish was not plasma treated and the solution was not left for evaporation.
FNDs with a nominal size of 70 nm were used for sensing experiments as they are preferably used in our most recent cell experiments [Niora2024]. The FNDs (Sigma-Aldrich 1003626896) were prepared following the same protocol. A suspension of FNDs was created by diluting 1 µL of the commercial stock solution in 1000 µL of deionized water, followed by sonication for 10 minutes to ensure a homogeneous dispersion. Approximately 50 µL of the suspension was deposited in a sterile Petri dish that was freshly unwrapped.
II.2 Optical Setup
The dried FND sample was placed inside a mini-incubator (pecon XS 2000) in an optical setup based on a customized Thorlabs OTKB/M modular optical tweezer system. Customization included a different objective and long-distance condenser lens. Vendors and other details are listed in Table 2 in the Appendix. The setup is constructed around an inverted microscope configuration, equipped with a high-resolution camera that capture wide-field bright-field images of the FND, and two lasers for the simultaneous illumination and manipulation of the FNDs. A green laser with a wavelength of 532 nm was utilized for the excitation of the NV centers, enabling the initialization and readout of their spin states. The second laser is operating in the near-infrared (NIR) at a wavelength of 1064 nm and used for FND trapping. This wavelength falls within the biological transparency window, which minimizes absorption and reduces the risk of significant heating effects in biological samples. The two lasers were aligned and their beams combined using a dichroic mirror, ensuring that they shared the same optical path before being focused onto the sample through the microscope objective, as detailed in Fig. 2a). By trapping an FND and then finetuning the position of the green laser to maximize the signal on the photon counter, we ensured the same focal point for both the green and NIR lasers. Additionally, acousto-optic deflectors were used to precisely control the timing and intensity of the laser beams by using Qudi software, enabling fine-tuned manipulation of the FNDs.
As illustrated in Fig. 2a), several modes of detection were installed: i) camera detection, ii) spectrometer detection and iii) photon count detection. A motorized flip-mirror ensures the ability to rapidly shift between camera detection and either spectrometer or photon count detection. The fluorescence spectrum of the FNDs is recorded and analyzed by a fiber coupled spectrometer while alternatively photon counts are measured with a fiber coupled avalanche photo detector. In addition, to measure the CW ODMR spectrum of FNDs, a signal generator connected to an amplifier is connected to a single-turn copper loop antenna (5 mm) placed approximately 1 mm above the sample. When the flip-mirror is positioned towards the camera, the latter captures high-resolution images of the FNDs illuminated during the experiments in real time. The combination of different detection methods thus made it possible to continuously monitor the position of the FNDs, determine whether they were isolated or in aggregates, and analyse their photoluminescence properties throughout the experimental process.
Figure 2: a) Illustration of the optical setup with the most important lasers, optical lenses, filters and equipment. Details of the components are specified in Table 2 in the appendix. b) Laser sequence for fluorescence relaxometry measurement. In green the 532 nm laser and in red the 1064 nm laser. Pulse-widths are $t_{\rm trap}=20\mu s$, $t_{ini}=35\mu s$, $t_{RO}=15\mu s$ while the variable waiting time $\tau$ allows to record the relaxometry at times spanning from $7.2\mu s$ to $720\mu s$.
II.3 FND Trapping
All trapping experiments were performed in a fluid chamber containing 50 µL of the FND aqueous suspension unless stated otherwise. The NIR laser was utilized for trapping the FNDs in the focused beam [Neuman2004, Gennerich2017, Jones2015]. For relaxometry, the temporal sequence of the green and trapping lasers, as illustrated in Fig. 2b), avoids temporal overlap between the two lasers while maintaining the FND stably trapped [Russell2018a]. Strictly speaking, this protocol does not constitute a conventional spin-relaxometry measurement. With an all-optical protocol, the method inherently involves potential inaccuracies due to photoionization induced by the green laser itself, thus affecting the NV- charge states. To address this limitation, measurement protocols incorporating microwave pulses explicitly designed to disentangle spin dynamics from photoionization effects have been developed. Examples include the microwave $\pi$-pulse method introduced by Ref. [Jarmola2023], as well as alternative sequences described in Ref. [Cardoso2023]. We hence will refer to the measured decay time as $T_{\rm relax}$ throughout the remainder of this paper. The green laser was pulsed with initialization time $t_{ini}$ of 35 µs, and read-out time $t_{RO}$ of 15 µs, for the readout of the NV center fluorescence, while the NIR laser was on to trap the FNDs in sequences with duration $t_{\rm trap}$ of 20 µs.
II.4 Sensing Experiments
For the experiments detecting paramagnetic species, a gadolinium solution (0.45 M in acetate) was gradually added to the FND’s environment. The concentrations were incrementally adjusted, and measurements were performed after each addition to evaluate the changes in the NV center properties.
For the pH sensing experiments, an initial 0.1 M NaOH(aq) solution at pH 13 was prepared. The pH was then progressively decreased by adding controlled volumes of a 1 M HCl solution. After each pH adjustment, the properties of the NV centers were measured while the FNDs were maintained in the optical trap at 60 mW.
For the temperature sensing experiments, the medium containing the FNDs was initially maintained at 37 C using an incubator. The temperature was then increased in steps of 1K, allowing the evaluation of temperature changes on the NV center properties. After each change of the set temperature of the incubator, we waited 10 minutes before the measurements were conducted.
For the sensing of paramagnetic species and pH, changes to the surrounding medium was involved. To ensure that the FND remained trapped despite disturbances in the medium, the CW NIR laser power was temporarily increased to 100 mW while liquid was added. Subsequently, the laser power was reduced back to 60 mW for measurements with the new conditions, and the same FND.
III Results
We begin by presenting the main results of our characterisation experiments, starting with the examination of the general fluorescence properties of NV-centers within the FNDs through photoluminescence (PL) spectroscopy. This foundational assessment provides an overview of the baseline optical characteristics and helps identify any immediate changes induced by the NIR laser irradiation. Following the PL spectrum analysis, we present the variations in fluorescence intensity as a function of optically detected magnetic resonance (ODMR) conditions. This step is crucial for understanding the impact of NIR irradiation on the magnetic resonance properties of the NV centers. Finally, we conduct fluorescence relaxometry to investigate the spin relaxation dynamics under different NIR laser parameters. PL spectrum measurement and ODMR measurement were conducted with CW NIR laser irradiation.
The experiments conducted under dry conditions primarily serve as a thermal control benchmark. This dry setup facilitates the understanding of thermal phenomena, such as localized heating effects, by providing a clear reference to compare with conditions involving FNDs suspended in aqueous environments. We acknowledge that the dry condition itself lacks direct practical application in sensing contexts.
For each set of experiments described in the subsections below, we exposed single FNDs to CW NIR laser at power levels of 0 mW, 30 mW, and 60 mW.
III.1 Photoluminescence Spectrum Analysis
As the NV neutral and negative charge states have distinct optical emission spectra, we conducted PL spectroscopy of FNDs at varying NIR power to investigate its impact. The FNDs were either dried on glass substrates or trapped in water to examine the influence of the surrounding medium. We record the PL spectrum in the range between 550 nm to 745 nm with the spectrometer for each of the different values of the CW NIR laser power.
Figure 3: Effect of CW NIR laser on dry FND emission spectrum. Optical emission spectrum of one FND, dry on glass, under CW NIR laser exposure at 0 mW (green), 30 mW (orange) and 60 mW (red) of 1064 nm laser power.
Upon CW NIR laser light exposure, our PL measurements revealed a significant drop in the total PL intensity. Notably, as a function of NIR laser power there was a slightly more pronounced reduction in the spectral intensity of NV- compared to NV0. For the fluorescence spectrum (like Fig. 3), the curves remain quite similar when water is added, although with less effect of increase in NIR laser intensity from 30mW to 60mW (data not shown) and the difference between 30 mW and 60 mW is even more reduced for FNDs that are trapped in water.
To quantify the changes in relative charge state emission intensities under CW NIR laser irradiation, we determine the ratio $\sigma_{0-}$ between NV0 and NV- emission intensities by integrating over the spectral ranges for NV0 (550–620 nm) and NV- (655–745 nm), respectively. Compared to the analysis used in Ref. [Sow2020], we applied a normalization factor 19/14 to account for the difference in the relative spectral ranges:
$\sigma_{0-}\equiv\left(\frac{\sum_{\lambda=550}^{620}I_{\lambda}}{\sum_{\lambda=655}^{745}I_{\lambda}}\right)\times\frac{19}{14}$ (1)
For FNDs trapped with NIR powers of 30 mW and 60 mW, we observe an averaged similar emission ratios $\sigma_{0-}$ of $0.48\pm 0.09$. In contrast, for FNDs dried on glass, we measure an emission ratio of $0.43\pm 0.11$ without NIR irradiation that increases to $0.57\pm 0.15$ for a NIR laser power of 60 mW NIR. All results are displayed in Fig. 6 parts a) and c).
III.2 Optically Detected Magnetic Resonance Measurements
In the next step, we elaborate on the influence of CW NIR laser light on the ODMR resonance frequency and contrast, revealing the impact of the NIR trapping laser on the NV electron spin properties.
The ODMR spectra were acquired in continuous-wave (CW) mode with constant intensities of both the 532 nm pump and 1064 nm trapping lasers and the microwave source, while sweeping the frequency of the microwave in steps of 0.3 MHz. Generally and for all FNDs, we observe a redshift in the resonance frequency with increasing CW NIR laser power, as illustrated in Fig. 4 by the spectra from one FND with 0 mW and 60 mW NIR power.
Figure 4: Effect of CW NIR laser power on ODMR spectrum of dry FND. ODMR spectra of one FND under CW NIR laser exposure at 0 mW (green) and 60 mW (red) of 1064 nm laser power. The resonance frequency shift and decrease in contrast with increasing CW NIR power is observed. The power of the 532 nm laser was 0.7 mW, measured before the objective lens, and the microwave amplifier was set to -2dB.
For FNDs dried on glass substrates, the resonance frequency shifted significantly to the lower frequencies. Taking 2.87 GHz as reference value for the zero field splitting, the resonance frequency shift increased in absolute value from $0.31\pm 0.54$ MHz at 0 mW to $5.83\pm 046$ MHz at 60 mW of CW NIR power. In contrast, FNDs trapped in water exhibited much smaller frequency shifts, with changes of $0.84\pm 0.21$ MHz at 30 mW and $1.58\pm 0.33$ MHz at 60 mW. Examples are shown in Fig. 4 and combined results displayed in Fig. 7. As before, for the similar experiment when water was added to the same FND that remained stuck on the glass slide, the ODMR spectra show a trend very similar to Fig. 4 although the difference between 30 mW and 60 mW of NIR power is less pronounced (data not shown).
[…]
III.3 Relaxometry Measurements
Last, we conducted relaxometry experiments under different NIR laser power, to extract values for the corresponding $T_{\rm relax}$. As above, FNDs were either dried on glass substrates or suspended in water to assess the influence of the surrounding medium on the relaxometry measurements with NIR irradiation. For these experiments, the all-optical pulsing sequence in Fig. 2b) was applied.
[…]
Figure 5: Effect of NIR laser power on $T_{\rm relax}$ relaxation times of dry FNDs. $T_{\rm relax}$ relaxometry measurements for one FND dried on glass under NIR laser powers of 0 mW (green) and 60 mW (red) at 1064 nm wavelength. Pulsing as in Fig 2b).
[…]
Overall, as showed in Fig. 6 and Fig. 7, the data suggest that the aqueous environment reduces the effect of NIR irradiation on ODMR-frequency shifts. Further, our results confirm that with specific NIR irradiation sequences and optimizing laser parameters, cf. Fig. 2b) and Ref. [Russell2018a], potential detrimental effects of NIR on $T_{\rm relax}$ relaxation can be significantly mitigated.
Figure 6: a) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs trapped inside water. b) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs trapped inside water. c) Correlation between NIR laser power, ODMR resonance frequency shift and $\sigma_{0-}$ fluorescence intensity ratio for FNDs dry on glass. d) Correlation between NIR laser power, ODMR resonance frequency shift and fluorescence relaxometry times for FNDs dry on glass. Data for 30 FNDs exposed to 0 mW (green), 30 mW (orange), and 60 mW (red) of 1064 nm NIR laser irradiation. CW NIR for ODMR and spectra, pulsed NIR as in Fig 2b) for fluorescence relaxometry. In general, we observe what appears to be a trend with differences between the three powers of NIR laser light for FNDs dry on glass, and and no or much less pronounced trend for FNDs in the optical trap in water.
Figure 7: Summary of experimental results, showing variation for individual nanodiamonds. (a) ODMR resonance frequency shift, relative to the expected resonance at 2.87 GHz, (b) fluorescence intensity ratio $\sigma_{0-}$, and (c) $T_{\rm relax}$ relaxometry times as functions of NIR laser power. Data for 5 random FNDs dry on glass (black stars) and trapped in water (gray circles) with 1064 nm NIR laser irradiation.
To study the significance of the observed effects, we performed an Analysis of Variance (ANOVA) test on the $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios. The ANOVA results are summarized in Table 1 (one-way Kruskal–Wallis test, $\alpha=0.05$).
Table 1: ANOVA test results for the effect of NIR laser power on FNDs dried on glass substrates, on glass after adding water and trapped in water. The p-values indicate the significance of differences observed across NIR power levels for $T_{\rm relax}$ relaxation times, ODMR frequency shifts, and $\sigma_{0-}$ ratios.
p-value
FNDs dried on glass
$T_{\rm relax}$ Relaxation Time <0.0001<0.0001
ODMR Frequency Shift <0.0001<0.0001
$\sigma_{0-}$ Ratio 0.0024
FNDs on glass with water
$T_{\rm relax}$ Relaxation Time $0.072$
ODMR Frequency Shift $0.008$
$\sigma_{0-}$ Ratio 0.201
FNDs Trapped in Water
$T_{\rm relax}$ Relaxation Time 0.221
ODMR Frequency Shift 0.073
$\sigma_{0-}$ Ratio 0.371
For FNDs dried on glass substrates, the very low p-values (p <0.01<0.01) for all three parameters confirm that the differences observed with varying NIR laser powers are statistically significant. This indicates that NIR irradiation has a considerable impact on the NV centers’ properties in air.
For FNDs on glass substrates after adding water, the mean p-values ($0.01<p<0.2$) for all three parameters confirm that the differences observed with varying NIR laser powers are not always statistically significant. In fact, the difference between no NIR laser and NIR laser is significant but the differences between 30 mW and 60 mW power of NIR laser are not significant.
In contrast, for FNDs trapped in water, the p-values are high (p >0.07>0.07), suggesting that the differences observed across different NIR power levels are not statistically significant. This reinforces the conclusion that the aqueous environment mitigates the effects of NIR irradiation on the NV centers.
III.4 Model of Charge State Dynamics
To understand and quantitatively model the effects of NIR laser illumination on the charge state dynamics and fluorescence properties of NV centers in FNDs, we investigate a comprehensive rate equation model [Qian2022], supplemented with insights from [Meirzada2018]. This model accounts for the various optical transitions, inter-system crossings, ionization, and recombination processes that occur when NV centers are simultaneously exposed to green and NIR laser excitation.
The NV center can exist in two charge states: the negatively charged NV- and the neutrally charged NV0. Each charge state has its own ground and excited electronic states, as well as inter-system crossing pathways to metastable quartet states. The model considers 8 states as illustrated in the schematic energy level diagram in Fig. 8.
The population dynamics of the NV centers are described by a set of coupled differential equations representing the rates of change of the population probabilities $P_{i}$ in each state [Qian2022]. The rates of change were adapted to our experimental situation with just two lasers, the green 532 nm and the NIR 1064 nm laser, following [Meirzada2018]. Values used in this work for the rates are provided in the appendix, Table 3.
The model is based on the following assumptions:
  • Steady-State Approximation: The populations reach a steady state within the simulation time, allowing us to analyze the equilibrium behavior.
  • Neglect of Higher-Order Processes: Multi-photon processes and higher excited states are neglected due to their low probabilities under the experimental conditions.
  • Constant Rate Coefficients: The rate coefficients are assumed to be constant and independent of the local temperature and environmental factors.
  • No External Fields: Effects of external magnetic or electric fields on the transition rates are not included.
Figure 8: Energy level diagram of the NV center, showing the ground states (P1, P2 and P7), excited states (P3, P4 and P8), quartet states (P5 and P6), and the transitions between them. Solid arrows represent radiative transitions, dotted arrows represent non-radiative transitions, and dashed arrows represent ionization and recombination processes.
The simulation results reveal that at low CW NIR laser powers, the population of NV- centers increases slightly, while the NV0 population decreases. This is attributed to enhanced recombination processes facilitated by the NIR photons, which convert NV0 centers back to NV-. Specifically, up to an CW NIR power of approximately 10 mW, the NV- population increases by about 8% compared to the case without CW NIR illumination.
However, as the CW NIR power increases beyond this threshold, the trend reverses. The NV- population begins to decrease, and the NV0 population increases. At a CW NIR power of 60 mW, the NV- population decreases by approximately 10% from its maximum value, while the NV0 population increases correspondingly. This inversion is due to the saturation of recombination processes and the dominance of ionization processes at higher CW NIR powers, which ionize NV- centers back to NV0.
The total fluorescence intensity from the NV centers decreases with increasing CW NIR power (Fig. 9). This decrease is primarily due to the reduced population of NV- centers and the lower fluorescence quantum yield of NV0 centers. The fluorescence from NV- centers dominates the emission under low CW NIR power, but as the CW NIR power increases, the contribution from NV0 centers becomes more significant and we observe a saturation in the fluorescence.
The $\sigma_{0-}$ ratio increases with CW NIR power. At low CW NIR powers, this ratio is low due to the predominance of NV- fluorescence. As the CW NIR power increases, the ratio rises slowly, indicating a relative increase in NV0 emission. This change in the fluorescence ratio affects the overall spectral characteristics and can influence the detection sensitivity in applications relying on NV- fluorescence.
The trends observed in the simulation align with the experimental results. The initial increase and subsequent decrease of the NV- population with increasing NIR power explain the non-monotonic behavior of fluorescence intensity (Fig. 9) and $\sigma_{0-}$ ratio (Fig. 6a,c)) observed experimentally.
Figure 9: Total fluorescence intensity from NV centers as a function of NIR laser power. (Black: Theory, Red: Experimental observations with errorbars, obtained as a statistical average over results from 10 different FNDs with $\sigma_{0-}$ within the range of $\sigma_{0-}=0.42\pm 0.04$.)
III.5 Temperature Simulation
To model the temperature increase induced by the NIR laser on FNDs, we simulated the heat transfer. The simulation accounts for the absorption of laser energy by the FND and the subsequent heat dissipation into the surrounding medium (water or air).
To quantify the temperature rise produced by the 1064 nm laser on a spherical fluorescent nanodiamond (radius $R=35\;\text{nm}$), we solve Fourier’s heat‑conduction equation with an internal absorption source inside the particle and pure conduction in the surrounding infinite medium (water or air):
$\rho_{d}c_{d}\frac{\partial T_{d}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{d}r^{2}\frac{\partial T_{d}}{\partial r}\right)+q_{\mathrm{abs}}(r),$ $0\leq r<R,$
$\rho_{m}c_{m}\frac{\partial T_{m}}{\partial t}$ $=\frac{1}{r^{2}}\frac{\partial}{\partial r}\!\left(k_{m}r^{2}\frac{\partial T_{m}}{\partial r}\right),$ $r>R,$
subject to the interface and far‑field conditions
$T_{d}(R,t)=T_{m}(R,t),$
$k_{d}\!\left.\frac{\partial T_{d}}{\partial r}\right|_{r=R}=k_{m}\!\left.\frac{\partial T_{m}}{\partial r}\right|_{r=R},\qquad\lim_{r\to\infty}T_{m}(r,t)=T_{\infty}.$
Here $\rho$, $c$, and $k$ are the density, specific heat, and thermal conductivity of the nanodiamond and the surrounding medium. The laser absorption is treated as a uniform volumetric heat source inside the FND,
$q_{\mathrm{abs}}(r)=\begin{cases}\dfrac{P_{\mathrm{abs}}}{\tfrac{4}{3}\pi R^{3}},&r<R,\\[6.0pt] 0,&r\geq R,\end{cases}$
where $P_{\mathrm{abs}}=\sigma_{\mathrm{abs}}I_{0}$ is the absorbed optical power for a beam of intensity $I_{0}$ and absorption cross‑section $\sigma_{\mathrm{abs}}$.
The simulation is based on the following assumptions:
  • Uniform Temperature Distribution: The FND is assumed to have a uniform temperature at any given time due to its small size.
  • Negligible Radiation Losses: Heat loss due to thermal radiation is neglected because it is insignificant compared to conduction at the temperatures considered.
  • Constant Thermal Properties: The thermal properties of the FND and the medium are assumed constant over the temperature range studied.
  • Steady-State Heat Transfer Coefficient: The convective heat transfer coefficient is considered constant and based on the thermal conductivity of the medium.
  • No Phase Change: The medium remains in the same phase (liquid for water, gas for air) throughout the simulation.
[…]
In Water: The temperature increase of the FND is minimal due to the high thermal conductivity and heat capacity of water. The simulation indicates that during the 500 ns NIR pulse:
  • The maximum temperature rise ($\Delta T_{\text{max}}$) is approximately 2 K above the ambient temperature.
  • The temperature increases steadily during the laser pulse and rapidly returns to ambient temperature within nanoseconds after the laser is turned off.
In Air: The temperature increase is more pronounced because air has lower thermal conductivity and heat capacity. The simulation results show:
  • The maximum temperature rise reaches approximately 41 K above the ambient temperature during the 500 ns NIR pulse.
  • The temperature increases during the laser pulse and decreases more slowly than in water after the laser is turned off.
The simulation results align with experimental observations where pulsed NIR illumination causes negligible thermal effects in aqueous environments, cf. Fig. 7. In experiments conducted in air, temperature increases contribute to shifts in ODMR frequency and should be accounted for. These shifts correspond to a 2 MHz shift in the ODMR resonance frequency which is what we experimentally observe.
A direct comparison of Fig. 6c)–d) with Fig. 6a)–b) shows that the environment (air versus water) is the controlling parameter. In air, the ODMR resonance shifts by $5.8\pm 0.5$ MHz and $T_{\rm relax}$ falls by $(44\pm 6)\%$ when the CW NIR power is raised from 0 to 60 mW, whereas in water the corresponding changes are only $1.6\pm 0.3$ MHz and $(2\pm 5)\%$, respectively. By contrast, the NV0/NV- fluorescence intensity ratio $\sigma_{0-}$ responds similarly to NIR power in both media (Fig. 6a,c; projection on the ordinate axis), confirming that charge-state kinetics are governed by photon flux rather than heat removal.
For completeness we note that NV- centres possess a weak intersystem transition between the metastable singlet states $^{1}E\,(p_{6})\!\rightarrow\!{}^{1}A_{1}\,(p_{5})$ with a zero-phonon line at $\lambda_{\mathrm{ZPL}}=1042$ nm and a room-temperature homogeneous width of only a few GHz ($\lesssim 10^{-4}$ nm) [Acosta2010, Rogers2015]. Our trapping beam at 1064 nm is therefore red-detuned by $\Delta\lambda=22$ nm, corresponding to $\Delta\nu\simeq 6$ THz ($\simeq$  25 meV), i.e. $\sim 10^{3}$ times the natural linewidth and well outside the vibrational sideband of the singlet manifold. Values for the measured absorption cross-sections decrease from $\sigma\!\approx\!2\times 10^{-19}$ cm2 on resonance to below $10^{-21}$ cm2 at this detuning [Acosta2010], so direct optical pumping of the singlet transition by the 1064 nm light is negligible. The effects reported in this work (PL quenching, ODMR shifts, $T_{\rm relax}$ variations) therefore arise from photothermal heating and photo-ionisation dynamics, not from resonant excitation of the 1042 nm line.
IV Sensing Capabilities of Trapped FNDs
As a final step in the comprehensive investigations, this section explores the potential of optically trapped FNDs with NV centers for sensing applications. Experiments focused on three key parameters: Gd3+ concentration, pH, and temperature. Measurements included ODMR, fluorescence relaxation, and optical spectrum analysis under variation of the three parameters mentioned above. Results are summarized in Fig. 10.
Figure 10: Variation in three sensing parameters when subject to controlled variation of paramagnetic species, pH, and temperature. a) ODMR resonance frequency shift VS Gd3+ concentration. b) ODMR resonance frequency shift VS pH value. c) ODMR resonance frequency shift (relative to the expected 2.87 GHz) VS Temperature increase (relative to 37 C). d) Fluorescence relaxometry time $T_{\rm relax}$ VS Gd3+ concentration. e) $T_{\rm relax}$ VS pH value. f) $T_{\rm relax}$ VS Temperature increase. g) $\sigma_{0-}$ fluorescence intensity ratio VS Gd3+ concentration. h) $\sigma_{0-}$ fluorescence intensity ratio VS pH value. i) $\sigma_{0-}$ fluorescence intensity ratio VS Temperature increase. The boxplot represent results for 10 FNDs that were trapped for each condition with 60 mW 1064 nm CW NIR laser. In each case, results for 0 M Gd3+, pH 13 and 0 K trapped were used as reference values for the individual FNDs. The red squares emphasize the optimal measurement method for each sensing parameter.
IV.1 pH Sensing
The impact of environmental pH on the properties of NV centers was evaluated using buffer solutions with pH values of 1, 7, and 13.
Figure 11: $\sigma_{0-}$ fluorescence intensity ratio during pH value changes. Measurement of the same FND while changing from pH13 to pH1 by adding HCl solution. Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • Fluorescence relaxation: A decrease in $T_{\rm relax}$ values was observed as pH increased, indicating a dependence on surface protonation dynamics (Fig. 10.h). This is consistent with reports that high pH conditions stabilize spin relaxation dynamics [Fujisaku2019, Rendler2017].
  • Optical Spectrum: A decrease in the $\sigma_{0-}$ ratio with increasing pH (Fig. 10.e and Fig. 11) suggests that alkaline conditions favor the NV- charge state. It may be attributed to the deprotonation of surface functional groups such as carboxyl (–COOH) and hydroxyl (–OH) groups present on the nanodiamond surface. Deprotonation of these groups under alkaline conditions generates negatively charged surface sites, thereby enhancing electron donation to nearby NV centers and favoring the stabilization of the NV- charge state [karaveli2016]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (Fig. 11) shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant changes in resonance frequency (Fig. 10.b) were detected across the tested pH range, indicating minimal impact of pH on electronic spin states.
The $\sigma_{0-}$ ratio’s dependence on pH underscores the importance of surface chemistry in modulating charge state dynamics, especially with FNDs in suspension.
To further explore the dependence of the NV charge state on pH, Fig. 11 presents the evolution of the $\sigma_{0-}$ ratio when varying pH from 13 to 1 for the same FND, both under trapping conditions and without NIR laser exposure. The trapped FNDs in suspension exhibit a significantly stronger modulation of $\sigma_{0-}$ compared to the same particle measured on glass. This enhanced responsiveness highlights the improved surface accessibility and chemical exchange in solution, reinforcing the utility of optically trapped FNDs for pH nanosensing. The pronounced increase in NV center fraction under alkaline conditions (lower $\sigma_{0-}$) confirms the surface deprotonation hypothesis, while acidic conditions suppress this stabilization. Overall, these results validate that $\sigma_{0-}$ serves as a reliable and sensitive optical proxy for probing local pH variations in biological or environmental settings.
IV.2 Paramagnetic Species Sensing
The sensitivity of NV centers to paramagnetic gadolinium ions (Gd3+) was investigated through measurements at 3 different values of the concentration.
Figure 12: $\sigma_{0-}$ fluorescence intensity ratio during gadolinium concentration changes. Measurement of the same FND while changing gadolinium concentration from 0M to 0.1M by adding a gadolinium solution (0.45 M in acetate). Red: FND trapped with 60 mW 1064 nm CW NIR laser, Green: FND on glass without NIR laser.
  • Fluorescence relaxation: A significant decrease in $T_{\rm relax}$ values was observed with increasing Gd3+ concentration (Fig. 10.g). This result aligns with the quenching effect of paramagnetic ions, which enhance spin-lattice relaxation by introducing magnetic noise to the local environment [Martinez2020, Rendler2017].
  • Optical Spectrum: A decrease in $\sigma_{0-}$ was observed (Fig. 10.d and Fig. 12). This can be attributed to charge state dynamics, where Gd3+ ions interact with the FND surface, altering the local electrochemical potential and promoting ionization of NV- to NV0 [Rondin2010]. The difference in ratio change between FNDs dry on glass without an NIR laser and FNDs trapped with a 60 mW CW IR laser (fig. 12) shows a clear advantage of using FNDs in suspension for sensing.
  • ODMR: No significant shifts or changes in resonance frequency were detected, suggesting limited interaction between Gd3+ ions and the electronic spin states of the NV centers under the tested conditions (Fig. 10.a).
These findings demonstrate the potential of NV centers for detecting paramagnetic species like Gd3+. The sensitivity saturates at higher concentrations due to the finite interaction volume and saturation of surface sites for Gd3+ binding. The fluorescence relaxometry provides a robust method for quantifying Gd3+ in biological and environmental contexts. However, the effect on charge state dynamics also imply that conclusions from fluorescence relaxometry in otherwise uncharacterized environments should be made with caution, and at best complemented with fluorescence spectrometry. For completeness, we observe that by adding Gd3+ to a concentration of 0.1 M, the pH value in the solution increases slightly (by of order 1-2 pH units) which is consistent with our previous observations on the effect of pH on the fluorescence of FNDs.
IV.3 Temperature Sensing
The temperature response of NV centers was evaluated under three conditions: FNDs on glass substrates with no NIR laser irradiation, trapped FNDs with no temperature increase, and trapped FNDs with temperature increases of +1 K and +2 K.
  • ODMR: A temperature-dependent shift in the resonance frequency was observed, with lower frequencies at higher temperatures (Fig. 10.c). This shift is attributed to the well-known temperature dependence of the NV center’s zero-field splitting parameter $D$, which decreases at a rate of approximately $74~\text{kHz/K}$ [Acosta2010].
  • Fluorescence relaxation: No significant changes were detected in $T_{\rm relax}$ values, indicating that spin-lattice relaxation is relatively insensitive to temperature variations within the tested range (Fig. 10.i).
  • Optical Spectrum: No significant changes in $\sigma_{0-}$ were observed with temperature, suggesting that charge state dynamics are minimally affected by thermal fluctuations (Fig. 10.f).
ODMR’s temperature sensitivity highlights its utility for high-resolution nanoscale thermometry. The absence of significant $T_{\rm relax}$ or optical spectral changes simplifies calibration, making ODMR a standalone method for temperature measurements
V Discussion
The shift in the ODMR resonance frequency with increasing NIR power, as shown in Fig. 4, Fig. 6, and Fig. 7a), suggests an increase in temperature within the FNDs. This is consistent with the known temperature dependence of the zero-field splitting parameter $D$ of the NV center, which decreases with increasing temperature at a rate of approximately 74-74 kHz/K [Acosta2010, Toyli2012]. However, the extent of the thermal effect is significantly influenced by the surrounding medium’s properties. Media with higher thermal conductivity and heat capacity, such as water, dissipate heat more efficiently than air, resulting in smaller temperature increases in the FNDs. This is supported by our temperature simulations, which show minimal temperature rise in water compared to air during NIR illumination. Consequently, FNDs trapped in water exhibit much smaller ODMR frequency shifts with increasing NIR power, as shown in Fig. 6 and Fig. 7.
In addition to thermal effects, the NIR laser alters the charge-state dynamics of the NV centers. Our simulations of charge-state dynamics indicate that at low NIR power levels, there is a slight increase in the NV- population due to enhanced recombination processes facilitated by NIR photons. However, as the NIR power increases beyond a threshold (approximately 10 mW with our parameters), the NV- population begins to decrease while the NV0 population increases. This inversion is attributed to the saturation of recombination processes and increased ionization rates at higher NIR powers.
The total fluorescence intensity decreases with increasing NIR power, as shown in Fig. 9. This decrease can be explained by the reduced NV- population and the lower fluorescence quantum yield of NV0 centers [Chapman2010]. The increase in $\sigma_{0-}$ with higher NIR power, depicted in Fig. 6 and Fig. 7, further confirms the shift in charge state distribution. Additionally, the increased NV0 background fluorescence dilutes the ODMR contrast since NV0 centers do not exhibit spin-dependent fluorescence variations. This is reflected in the decreased ODMR contrast observed in our measurements (Fig. 4). Charge-state fluctuations also introduce additional pathways for spin relaxation, contributing to the reduction in $T_{\rm relax}$ times [Aslam2013, Goudarzi2019].
Our results indicate that although both thermal effects and charge-state dynamics significantly impact the properties of the NV centers under NIR illumination, their relative contributions depend on the environmental conditions. In FNDs dried on glass substrates, the pronounced ODMR frequency shifts suggest that thermal effects dominate due to inefficient heat dissipation in air. In contrast, for FNDs trapped in water, the minimal temperature increase reduces thermal effects, making charge-state dynamics more prominent in influencing the NV centers’ behavior. The increase in the $\sigma_{0-}$ ratio and the decrease of fluorescence relaxometry times with NIR power without significant ODMR frequency shifts (Fig. 6) indicate that charge-state transitions are the primary mechanism affecting the fluorescence and spin properties in aqueous environments.
While thermal effects and charge-state dynamics provide a comprehensive explanation for our observations, alternative mechanisms should be considered. Multi-photon absorption processes could directly excite the NV centers without significant heating, yet, the relationship between ODMR frequency shift and NIR power (Fig. 6) is contrary to this explanation. Moreover, the significant differences in the effects of NIR irradiation between FNDs in air and water emphasize the critical role of the surrounding medium. The statistical analysis using ANOVA (Table 1) confirms that the differences observed in FNDs dried on glass are statistically significant, while those in FNDs trapped in water are not.
In our opinion, the data presented in this work establish a clear hierarchy of influences: (i) the thermal conductivity and heat capacity of the surrounding medium dictate the magnitude of the NIR-induced temperature rise and therefore the thermally induced shifts in ODMR frequency and the shortening of $T_{\rm relax}$; (ii) the NIR photon flux sets the rate of photo-ionisation and recombination, which in turn controls the NV0/NV- balance and the fluorescence contrast. Optical trapping itself plays a mainly mechanical role-holding a single FND at a fixed position—while potential effects on spin and charge-state dynamics depend on the wavelength of the trapping laser.
This distinction is important for two reasons. First, it tells experimentalists which knob to turn: to suppress thermal artefacts one changes the medium (e.g. buffer composition, flow, heat sinking) rather than the trap geometry and wavelength. Second, it generalises our findings to intracellular sensing, where the cytoplasm’s thermal properties resemble those of water: the modest $\lvert\Delta f\rvert<2$ MHz and unchanged $T_{\rm relax}$ observed in vitro set an upper bound on in-cell NIR heating, validating the use of trapped FNDs for quantitative biosensing.
The results presented in Section IV confirm that optical trapping of FNDs provide a valuable add-on that allows to enhance the application potential of biosensing with NV-centers in nanodiamonds. In particular, Fig. 10a-c) demonstrates that the frequency shift in ODMR is a good stand-alone measure for a temperature shift also for optically trapped FNDs, while parts f) and i) illustrate that changes in the surrounding temperature within a range that is relevant for biological systems does not appear to alter the readings provided by the fluorescence intensity ratio $\sigma_{0-}$ or the relaxation time, $T_{\rm relax}$.
VI Conclusion
This work investigates the effects of near-infrared (NIR) laser irradiation on fluorescent nanodiamonds (FNDs) containing nitrogen-vacancy (NV) centers, offering a comprehensive understanding of the interplay between thermal effects, charge-state dynamics, and their combined influence on sensing capabilities. By employing an integrated experimental approach combining photoluminescence spectroscopy, optically detected magnetic resonance (ODMR), and fluorescence relaxometry, this study elucidates how varying NIR laser powers impact the fluorescence, spin properties, and stability of NV centers in diverse environmental conditions.
A key finding is the dual role of NIR irradiation: in dry environments, thermal effects dominate, resulting in significant shifts in ODMR frequencies and reductions in relaxation times due to localized heating. In contrast, for FNDs that are fully submerged in aqueous environments, the surrounding water appears to mitigate these thermal effects through efficient heat dissipation, allowing charge state dynamics to take precedence. The increased $\sigma_{0-}$ ratio and its implications for sensing applications underscore the critical role of the surrounding medium in tailoring NV center behavior. The study also demonstrates the sensitivity of optically trapped FNDs to pH variations, temperature fluctuations, and paramagnetic species, establishing these nanodiamonds as versatile tools for nanoscale sensing.
The findings indicate that optimizing NIR laser parameters, including power levels and pulsing sequences, can significantly enhance sensor performance while minimizing adverse effects such as fluorescence quenching and spin relaxation. Additionally, the demonstrated ability to detect and differentiate environmental factors such as pH and gadolinium ions highlights the potential of NV centers for applications in biological sensing and environmental monitoring.
However, this study also reveals critical uncertainties and challenges that must be addressed to fully exploit the potential of NV-based FNDs. Long-term stability under extended NIR irradiation remains an open question, particularly regarding potential degradation of fluorescence and spin properties. The effects of complex biological environments, including interactions with proteins, ions, and other cellular components, are yet to be fully understood. The role of surface functionalization in enhancing NV center stability and sensing accuracy requires further exploration, as does the potential for alternative NIR wavelengths or advanced laser modulation techniques to improve performance.
Theoretical modeling, while providing valuable insights into charge state dynamics and thermal effects, must be expanded to include the complexities of real-world applications. Future work could explore the integration of complementary sensing techniques, such as fluorescence lifetime imaging or advanced magnetic resonance methods, to provide a a real-time sensing protocol inside cells by using optically trap FNDs in the cellular medium. Furthermore, efforts to miniaturize and automate the experimental setup, enabling in vivo applications, would be a critical step in translating this research into practical diagnostic tools.
acknowledgments
This work was supported by the Independent Research Fund Denmark (grant no. 0135-00142B) and the Novo Nordisk Foundation (grant no. NNF20OC0061673). We acknowledge discussions with Maabur Sow and Fedor Jelezko, supported by ERC through HyperQ Project (SyG 856432). We further acknowledge insights provided through the student project work of Kristian Lambertsen and Nicolai Kongstad, as well as Suyash Amzare, including an original sketch of the experimental setup providing the basis for Fig. 2.
appendix
We specify in Table 2 all part numbers in the optical setup, Fig. 2a).
Table 2: Optical setup component
Ref. Number Component Model
1 IR laser Cobolt - Rumba 3000
2 IR polarizer ThorLabs Multi-order polarizer
3 IR beam dump ThorLabs BTC30
4 IR PBS ThorLabs CCM1-PBS253/M
5 Lens ThorLabs LA1708-B-ML
6 Acousto-Optic Deflector AA Optoelectronics MT200 - A0.5 - 1064nm
7 Iris Aperture ThorLabs ID12Z/M
8 IR disk ThorLabs VRC4CPT
9 Beam expander
10 Right-angle prism ThorLabs DFM1/M
11 Green laser Verdi G-Series - Coherent
12 Optical isolator EOT 04-532-00113
13 Green polarizer ThorLabs Multi-order polarizer
14 Green PBS ThorLabs VA5-532/M
15 Lens ThorLabs LA1131 - A - ML
16 Acousto-Optic Deflector Brimrose TEM-110-25-530
17 Iris Aperture ThorLabs SM1D12D
18 Lens ThorLabs LA1131 - A - M
19 Iris Aperture ThorLabs P30D
20 Lens ThorLabs LA1509 - A
21 Right-angle prism ThorLabs DFM1/M
22 Lens ThorLabs LA1484 - A
23 Filters Bightline Basic 460/60 - ThorLabs MF630-69 Emission filter
24 Camera Teledyne Prime 95B
25 Filter ThorLabs NE20A
26 Filters Bightline Basic FF01-630/52 - ThorLabs FELH 0550
27 Objective lens Olympus PLN 10X
28 Single Photo counting module Excelitas SPCM-AQRM-13-FC
29 Iris aperture ThorLabs SM1D12D
30 Objective lens Nikon Plan Apo 100X
31 Microwave Antenna Copper wire
32 Microwave Generator Agilent N5181A
33 Condensor Mitutuyo Plan Apo 100X
34 Quadrant photoelectric detector ThorLabs PDQ80A
35 White lamp
36 Spectrometer Ocean Optics QE Pro
In literature, a varying range of values for the rate constants used in the numerical model corresponding to Fig. 8 exist. We applied the relevant values from Ref. [Qian2022] and combined them with parameters from [Meirzada2018] for the four ionization and recombination rates as these authors use the same laser wavelengths as us in their protocol. Below, in Table 3, we provide the values used in the simulations of Section III.4.
Table 3: Parameters used in the simulation
Parameter Value Units Description
$K_{\mathrm{NV}^{-}}^{(e)}$ 27 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ NV- excitation rate
$K_{\mathrm{NV}^{0}}^{(e)}$ 18 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ NV0 excitation rate
$\tau_{\mathrm{NV}^{-}}$ $13\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (13 ns)
$\tau_{\mathrm{NV}^{0}}$ $19\times 10^{-3}$ $\mu\mathrm{s}$ Excited-state lifetime (19 ns)
$K_{\mathrm{NV}^{-}}^{(f)}$ $1/\tau_{\mathrm{NV}^{-}}$ $(\mu\mathrm{s})^{-1}$ NV- fluorescence rate
$K_{\mathrm{NV}^{0}}^{(f)}$ $1/\tau_{\mathrm{NV}^{0}}$ $(\mu\mathrm{s})^{-1}$ NV0 fluorescence rate
$K_{35}$ 7.9 $(\mu\mathrm{s})^{-1}$ NV- ES ms=0 $\to$ singlet
[…]
$K_{56}$ 1000 $(\mu\mathrm{s})^{-1}$ Singlet ES $\to$ GS
$K_{61}$ 6.5 $(\mu\mathrm{s})^{-1}$ Singlet GS $\to$ NV- GS ms=0
[…]
$K_{iG}$ 852 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Ionization rate (green laser)
$K_{rG}$ 134 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Recombination rate (green laser)
$K_{iIR}$ 1.2 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Ionization rate (IR laser)
$K_{rIR}$ 3.17 $(\mu\mathrm{s})^{-1}\mathrm{/mW}$ Recombination rate (IR laser)
laser_green_power 0.7 mW Green laser power